# Gait Recognition in Large-scale Free Environment via Single LiDAR

### v2

@@ Introduction
Figure 3: The pipeline of our method. We extract dense body structure information from range views, and undistorted geometric and motion features via motion-aware feature embedding (MAFE) from point clouds. Then, adaptive cross-representation mapping module (ACM) is applied to fuse two-representation features at different levels hierarchically. Lastly, the gait-saliency feature enhancement (GSFE) module is leveraged to highlight more gait-informative features for final identification.
Problem Definition
@@
Given a LiDAR-based gait recognition dataset $\mathcal{P}=\{\mathcal{P}_{i}^{j}\mid i=1,2\ldots,N;j=1,2,\ldots,m_{i}\}$ with N individuals and $m_{i}$ sequences for each individual $d_{i}$. Each sequence $\mathcal{P}_{i}^{j}$ consists of fixed $T$ frames of raw LiDAR point clouds $Y=\left\{\mathbf{y}_{t}\right\}_{t=1}^{T}$ and projected LiDAR range views $X=\left\{\mathbf{x}_{t}\right\}_{t=1}^{T}$, where $\mathbf{y}_{t}\in\mathbb{R}^{N\times 3}$ is the $t_{th}$ frame of point clouds and $\mathbf{x}_{t}\in\mathbb{R}^{H\times W}$ is the $t_{th}$ frame of range views. For a given LiDAR point cloud of each subject, we convert each point ${p}_{i}=(x,y,z)$ to spherical coordinates and finally to 2D pixel coordinates, using the following projection function (Kong and etc. 2023):
@@
$\left(\begin{array}[]{l}u_{n}\\ v_{n}\end{array}\right)=\left(\begin{array}[]{c}\frac{1}{2}\left[1-\arctan(y,x)\pi^{-1}\right]w\\ {\left[1-\left(\arcsin\left(zr^{-1}\right)+\mathrm{f}_{\mathrm{up}}\right)\mathrm{f}^{-1}\right]h}\end{array}\right),$ (1)
@@
where $\left(u_{n},v_{n}\right)$ denotes range view coordinates, $\left(h,w\right)$ is the height and width of the desired range view predefined by the inherent parameters of the LiDAR. $\mathrm{f}=\mathrm{f}_{\mathrm{up}}+\mathrm{f}_{\text{down }}$ is the vertical field-of-view of the sensor, and $r=\left\|{p}_{i}\right\|_{2}$ is the range of each point. Then, we crop and resize the range view with each subject to a resolution of $64\times 64$. We aim to learn a network $N_{\theta}(\cdot)$ that can map the inputs to feature embedding $\mathcal{F}_{i}^{j}$ to represent the corresponding individual $d_{i}$.
Overview
@@
We propose a hierarchical multi-representation feature interaction network (HMRNet), a novel point-range gait recognition solution by taking advantage of LiDAR-projected range views and raw point clouds. Our pipeline’s overview is shown in Figure 3, taking a sequence of range viewss and point clouds as input to identify individuals based on their gait. There are three main procedures in our network, including hierarchical adaptive cross-representation mapping (H-ACM), motion-aware feature embedding(MAFE), and gait-saliency feature enhancement(GSFE). We employ ResNet-like CNNs (Fan and etc 2023) and MLPs (Ma, Qin, and etc 2022) to extract multi-resolution features from range views and point clouds, and fuse valuable geometric and dynamic information hierarchically. Notably, we leverage motion-aware feature embedding to explicitly model gait-related motion information from point clouds. After hierarchical multi-representation feature interaction, temporal pooling and horizontal pyramid pooling (HPP) are utilized following (Fan and etc 2023), to gather comprehensive fusion features. Before final identification, we employ gait-saliency feature enhancement module to highlight gait-informative features, benefiting from the channel-wise attention mechanism.
Hierarchical Adaptive Cross-representation Mapping
@@
$F_{attention}=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V.$ (2)
The final fusion feature $F_{fusion}$ is acquired through the feed-forward network (FFN) and layer normalization in transformer(Vaswani, Shazeer, and etc 2017) by
$F_{fusion}=LN(F_{attention}+FFN(F_{attention})).$ (3)
@@
Table 2: Comparison with SOTA methods of gait recognition on FreeGait and SUSTech1K.
@@
Silhouette GaitSet (Chao and etc 2019) 57.13 71.87 64.01 65.22 84.91 74.26
GaitPart (Fan, Peng, and etc. 2020) 52.26 64.95 58.54 59.29 80.79 69.18
GLN (Hou, Cao, and etc 2020) 52.21 68.69 60.01 65.78 84.76 74.49
GaitBase (Fan and etc 2023) 62.64 75.30 68.57 77.50 90.22 83.44
Silhouette&SMPL SMPLGait (Zheng and etc. 2022) 53.39 69.25 60.93 66.34 85.08 74.95
Silhouette&Key Point SMPLGait (Zheng and etc. 2022) 57.97 73.02 65.12 69.75 86.68 77.60
3D Key Point GaitGraph (Teepe and etc 2021) 14.69 27.81 21.74 2.09 6.59 5.27
Point Cloud PointNet (Qi and etc 2017) 42.48 62.48 52.05 37.31 65.01 50.11
PointMLP (Ma, Qin, and etc 2022) 57.61 76.68 66.30 68.86 90.32 78.55
LiDARGait (Shen et al. 2023) 74.15 88.75 80.66 86.81 95.98 91.08
@@
Motion-aware Feature Embedding
Body part movements across successive frames are essential for effective gait recognition. However, existing LiDAR-based gait recognition methods rarely model motion information explicitly. By incorporating 3D raw point clouds that retain comprehensive pedestrian motion information, we can explicitly model gait-related dynamic information through point-wise flow. In each stage of point branch, imagine each anchor point $p^{t}_{i}$ at frame $t$ after farthest point sampling(FPS), $p^{t}_{i}$ is represented by its Euclidean coordinates $\mathbf{x}_{i}^{t}\in\mathbb{R}^{3}$ and a feature vector $\mathbf{f}_{i}^{t}\in\mathbb{R}^{c}$ from point encoder with MLPs (Ma, Qin, and etc 2022). We learn to aggregate its local geometric and dynamic motion feature by the neighboring $k$ points ($k=16$) of $p^{t}_{i}$ in the same frame $\mathcal{N}(p_{i}^{t})$ and nearby frame $\mathcal{N^{\prime}}(p_{i}^{t-1})$. As shown in Figure 4, for each pair $(q_{j}^{t},q_{j}^{t-1})$ in $\mathcal{N}(p_{i}^{t})$ and $\mathcal{N^{\prime}}(p_{i}^{t-1})$ respectively, we pass the difference of their origin Euclidean coordinates into MLPs to obtain the motion-aware feature embedding ${m}_{j}^{t}\in\mathbb{R}^{c}$ in hidden states of neighboring point $q^{t}_{j}$. Then, we get the local geometric feature with motion information for neighboring points by element-wise addition of their geometric feature ${f}_{j}^{t}$ and motion embedding ${m}_{j}^{t}$. Finally, we aggregate them by element-wise pooling and update the feature vector for every anchor point $p^{t}_{i}$. The above operations can be formulated as:
$h(p_{i}^{t})=\underset{q_{j}^{t}\in\mathcal{N}(p_{i}^{t})}{MAX}\{(f_{j}^{t}+\ \zeta({x}_{j}^{t}-{x}_{j}^{t-1})\},$ (4)
@@
Gait-saliency Feature Enhancement
For the hierarchical fusion features, we use temporal pooling and Horizontal Pyramid Pooling (HPP)(Fan and etc 2023), producing a comprehensive feature vector with $p$ ($p=16$) strips $f_{s}\in\mathbb{R}^{c}$ ($c=512$).
Figure 4: The procedure of point-wise flow to learn motion relation between two adjacent frames.
Believing that different channel-wise feature maps represent various gait-related attributes with differing importance for gait recognition. To leverage this insight, we introduce a gait-saliency feature enhancement module that utilizes a channel-wise attention mechanism (Hu, Shen, and etc 2018). Our gait-saliency feature enhancement module can adaptively recalibrate channel-wise feature responses by explicitly modeling interdependencies between channels, and reweighting the input feature map to emphasize informative information. Consequently, our network efficiently highlights significant gait-related features while reducing the impact of redundant features, leading to more accurate gait recognition. Lastly, we combine gait-saliency enhancement features of various strips and obtain the final gait feature $f_{\beta}\in\mathbb{R}^{c}$ ($c=256$) for further loss computation through a feature head with multi-FC layers.
Experiments
Table 3: Evaluation with different attributes on SUSTech1K valid + test set. We compare our method with silhouette-based SOTA method GaitBase, LiDAR-based SOTA method PointMLP and LiDARGait.
@@
Silhouette GaitBase (Fan and etc 2023) 83.09 79.34 50.95 76.98 77.34 77.31 83.46 26.65 77.50 90.22
Point Cloud PointMLP (Ma, Qin, and etc 2022) 76.03 71.91 57.09 68.08 58.29 63.28 79.25 70.75 68.86 90.32
LiDARGait (Shen et al. 2023) 91.91 88.61 75.27 88.99 67.55 81.19 94.73 90.04 86.81 95.98
@@
Overview
In this section, we outline our evaluation, training, and inference details, followed by comprehensive comparison results. Ablation studies highlight our network’s effectiveness, and we also analyze its robustness under cross-views, low-light conditions and various sequence length.
Evaluation Protocol
@@
For FreeGait, we resize the silhouettes to $128\times 88$, similar to Gait3D (Zheng and etc. 2022). Each LiDAR point cloud frame is resampled to $N=256$ points via farthest point sampling(FPS), normalized by setting its center as the origin while preserving original orientations. Range views are cropped and resized to $64\times 64$, following LiDARGait (Shen et al. 2023). For SUSTech1K (Shen et al. 2023), we resample the point clouds to 512 points using FPS, as they have an average of 800 points compared to FreeGait’s 300 points. Other input settings remain consistent with the paper. All methods, except LiDARGait and our HMRNet, are trained with Adam optimizer with a weight decay of 0.0005 and an initial learning rate of 0.001 on the two benchmarks. We employ multi-step learning rate reduction by a factor of 0.1 at the 10,000th and 30,000th iterations, with a total of 50,000 iterations. For LiDARGait and our HMRNet, we maintain the same training procedure on both benchmarks. During the training stage, each input sequence consists of 10 frames, approximating the average length of a human gait cycle. In the training stage, the loss can be formulated as follows:
$L_{cls}=-\sum_{p=1}^{S}\sum_{c=1}^{C}gt_{s,c}\log\left(\operatorname{softmax}\left(f_{{\alpha}_{s}}\right)\right)_{c},$ (5)
where $gt_{s,c}$ indicates the identity information of the $s_{th}$ strip, which equals 0 or 1. $L_{tri}$ is used to optimize the inter-class and intra-class distance, which is computed by
$L_{tri}=\left[D\left(f_{\beta}^{A_{i}},f_{\beta}^{A_{j}}\right)-D\left(f_{\beta}^{A_{i}},f_{\beta}^{B_{k}}\right)+m\right]_{+},$ (6)
where $A_{i}$ and $A_{j}$ are samples from the same class $A$, while $B_{k}$ denotes the sample from another class. $D(d_{i},d_{j})$ is the Euclidean distance between $d_{i}$ and $d_{j}$ and $m$ is the margin of the triplet loss. The operation $[\gamma]_{+}$ equals to $max(\gamma,0)$. The overall loss function can be formulated as
$L=\alpha L_{tri}+\beta L_{cls}.$ (7)
We set the batch size to $(p=16,k=4)$ on our FreeGait and $(p=8,k=8)$ on SUSTech1K (Shen et al. 2023), where $p$ and $k$ denote the number of subjects and their corresponding training samples, respectively.
Comparison with SOTA Methods
@@
Table 2 compares our method with state-of-the-art (SOTA) gait recognition techniques on the FreeGait and SUSTech1K datasets. Our approach outperforms exisiting sota methods LiDARGait (Shen et al. 2023), showing improvements of $\mathbf{6.61\%}$ and $\mathbf{3.42\%}$ in Rank-1 accuracy for FreeGait and SUSTech1K, respectively. These achievements stem from our effective combination of 3D geometry and dynamic motion gait features. Silhouette-based methods suffer from distorted body shapes in view-dependent images, impacting performance. Though 3D SMPL and KP of FreeGait and SUSTech1K are based on the SOTA LiDAR-based mocap method (Zhao, Ren, and etc 2022), their accuracy is yet to be validated. We evaluated point-based gait recognition using PointNet(Qi and etc 2017) and PointMLP(Ma, Qin, and etc 2022) extractors, followed by a recognition head. These point-based methods grapple with sparse representation especially in distant scenes, but PointMLP(Ma, Qin, and etc 2022), leveraging local feature grouping, showcases the promise in point-based gait recognition. LiDARGait (Shen et al. 2023) mitigates point sparsity and disorder by converting them into a dense range-view representation, achieving notable results. However, this conversion can compromise the original rich 3D geometric and dynamic motion data from raw point clouds. Conversely, our HMRNet extracts dense body structures from range views and accurately captures the raw point clouds’ geometry and motion. It hierarchically processes both local fine-grained and global semantic features, benefiting gait feature extraction.
@@
Figure 5: The exemplar range views on SUSTech1K (Shen et al. 2023) in four different attributes with serious appearance variance between gallery set and probe set.
Ablation for Network Design.
@@
Table 4: Ablation studies for network modules on FreeGait.
@@
Table 5: Rank-1 accuracy under cross views and low-light conditions on FreeGait dataset.
@@
Silhouette GaitBase (Fan and etc 2023) 30.31 30.83
Point Cloud LiDARGait (Shen et al. 2023) 59.62 73.33
@@
More Analysis of Robustness.
@@
Figure 6: Average rank-1 accuracy with increasing testing frames on FreeGait dataset.

### v3

@@ 4. Methodology
Figure 4. The pipeline of our method. We extract dense body structure information from range views, and undistorted geometric and motion features via motion-aware feature embedding (MAFE) from point clouds. Then, adaptive cross-representation mapping module (ACM) is applied to fuse two-representation features at different levels hierarchically. Lastly, the gait-saliency feature enhancement (GSFE) module is leveraged to highlight more gait-informative features for final identification.
4.1. Problem Definition
@@
Given a LiDAR-based gait recognition dataset $\mathcal{P}=\{\mathcal{P}_{i}^{j}\mid i=1,2\ldots,D;j=1,2,\ldots,m_{i}\}$ with D individuals and $m_{i}$ sequences for each individual $d_{i}$. Each sequence $\mathcal{P}_{i}^{j}$ consists of fixed $T$ frames of raw LiDAR point clouds $Y=\left\{\mathbf{y}_{t}\right\}_{t=1}^{T}$ and projected LiDAR range views $X=\left\{\mathbf{x}_{t}\right\}_{t=1}^{T}$, where $\mathbf{y}_{t}\in\mathbb{R}^{N\times 3}$ is the $t_{th}$ frame of point clouds and $\mathbf{x}_{t}\in\mathbb{R}^{H\times W}$ is the $t_{th}$ frame of range views. For a given LiDAR point cloud of each subject, we convert each point ${p}_{i}=(x,y,z)$ to spherical coordinates and finally to 2D pixel coordinates, using the following projection function (Kong and etc., 2023):
@@
(1) $\left(\begin{array}[]{l}u_{n}\\ v_{n}\end{array}\right)=\left(\begin{array}[]{c}\frac{1}{2}\left[1-\arctan(y,x)\pi^{-1}\right]w\\ {\left[1-\left(\arcsin\left(zr^{-1}\right)+\mathrm{f}_{\mathrm{up}}\right)\mathrm{f}^{-1}\right]h}\end{array}\right),$
@@
where $\left(u_{n},v_{n}\right)$ denotes range view coordinates, $\left(h,w\right)$ is the height and width of the desired range view predefined by the inherent parameters of the LiDAR. $\mathrm{f}=\mathrm{f}_{\mathrm{up}}+\mathrm{f}_{\text{down }}$ is the vertical field-of-view of the sensor, and $r=\left\|{p}_{i}\right\|_{2}$ is the range of each point. Then, we crop and resize the range view with each subject to a resolution of $64\times 64$. We aim to learn a network $N_{\theta}(\cdot)$ that can map the inputs to feature embedding $\mathcal{F}_{i}^{j}$ to represent corresponding individual $d_{i}$.
4.2. Overview
@@
We propose a hierarchical multi-representation feature interaction network (HMRNet), a novel point-range gait recognition solution by taking advantage of LiDAR-projected range views and raw point clouds. Our pipeline’s overview is shown in Figure 4, taking a sequence of range views and point clouds as input to identify individuals based on their gait. There are three main procedures in our network, including hierarchical adaptive cross-representation mapping (H-ACM), motion-aware feature embedding (MAFE), and gait-saliency feature enhancement (GSFE). We employ ResNet-like CNNs (Fan and etc, 2023) and MLPs (Ma et al., 2022) to extract multi-resolution features from range views and point clouds, and fuse valuable geometric and dynamic information hierarchically. Notably, we leverage motion-aware feature embedding to explicitly model gait-related motion information from point clouds. After hierarchical multi-representation feature interaction, temporal pooling and horizontal pyramid pooling (HPP) are utilized following (Fan and etc, 2023), to gather comprehensive fusion features. Before final identification, we employ gait-saliency feature enhancement module to highlight gait-informative features, benefiting from the channel-wise attention mechanism.
4.3. Hierarchical Adaptive Cross-representation Mapping
@@
(2) $F_{attention}=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V.$
The final fusion feature $F_{fusion}$ is acquired through the feed-forward network (FFN) and layer normalization in transformer(Vaswani et al., 2017) by
(3) $F_{fusion}=LN(F_{attention}+FFN(F_{attention})).$
@@
Table 2. Comparison with SOTA methods of gait recognition on FreeGait and SUSTech1K.
@@
Silhouette GaitSet (Chao and etc, 2019) 57.13 71.87 64.01 65.22 84.91 74.26
GaitPart (Fan et al., 2020) 52.26 64.95 58.54 59.29 80.79 69.18
GLN (Hou et al., 2020) 52.21 68.69 60.01 65.78 84.76 74.49
GaitBase (Fan and etc, 2023) 62.64 75.30 68.57 77.50 90.22 83.44
Silhouette&SMPL SMPLGait (Zheng and etc., 2022) 53.39 69.25 60.93 66.34 85.08 74.95
Silhouette&Key Point SMPLGait (Zheng and etc., 2022) 57.97 73.02 65.12 69.75 86.68 77.60
3D Key Point GaitGraph (Teepe and etc, 2021) 14.69 27.81 21.74 2.09 6.59 5.27
Point Cloud PointNet (Qi and etc, 2017) 42.48 62.48 52.05 37.31 65.01 50.11
PointMLP (Ma et al., 2022) 57.61 76.68 66.30 68.86 90.32 78.55
LiDARGait (Shen et al., 2023) 74.15 88.75 80.66 86.81 95.98 91.08
@@
4.4. Motion-aware Feature Embedding
Body part movements across successive frames are essential for effective gait recognition. However, existing LiDAR-based gait recognition methods rarely model motion information explicitly. By incorporating 3D raw point clouds that retain comprehensive pedestrian motion information, we can explicitly model gait-related dynamic information through point-wise flow. In each stage of point branch, imagine each anchor point $p^{t}_{i}$ at frame $t$ after farthest point sampling(FPS), $p^{t}_{i}$ is represented by its Euclidean coordinates $\mathbf{x}_{i}^{t}\in\mathbb{R}^{3}$ and a feature vector $\mathbf{f}_{i}^{t}\in\mathbb{R}^{c}$ from point encoder with MLPs (Ma et al., 2022). We learn to aggregate its local geometric and dynamic motion feature by the neighboring $k$ points ($k=16$) of $p^{t}_{i}$ in the same frame $\mathcal{N}(p_{i}^{t})$ and nearby frame $\mathcal{N^{\prime}}(p_{i}^{t-1})$. As shown in Figure 5, for each pair $(q_{j}^{t},q_{j}^{t-1})$ in $\mathcal{N}(p_{i}^{t})$ and $\mathcal{N^{\prime}}(p_{i}^{t-1})$ respectively, we pass the difference of their origin Euclidean coordinates into MLPs to obtain the motion-aware feature embedding ${m}_{j}^{t}\in\mathbb{R}^{c}$ in hidden states of neighboring point $q^{t}_{j}$. Then, we get the local geometric feature with motion information for neighboring points by element-wise addition of their geometric feature ${f}_{j}^{t}$ and motion embedding ${m}_{j}^{t}$. Finally, we aggregate them by element-wise pooling and update the feature vector for every anchor point $p^{t}_{i}$. The above operations can be formulated as:
(4) $h(p_{i}^{t})=\underset{q_{j}^{t}\in\mathcal{N}(p_{i}^{t})}{MAX}\{(f_{j}^{t}+\ \zeta({x}_{j}^{t}-{x}_{j}^{t-1})\},$
@@ 4.4. Motion-aware Feature Embedding
4.5. Gait-saliency Feature Enhancement
For the hierarchical fusion features, we use temporal pooling and Horizontal Pyramid Pooling (HPP)(Fan and etc, 2023), producing a comprehensive feature vector with $p$ ($p=16$) strips $f_{s}\in\mathbb{R}^{c}$ ($c=512$).
Figure 5. The procedure of point-wise flow to learn motion relation between two adjacent frames.
Believing that different channel-wise feature maps represent various gait-related attributes with differing importance for gait recognition. To leverage this insight, we introduce a gait-saliency feature enhancement module that utilizes a channel-wise attention mechanism (Hu et al., 2018). Our gait-saliency feature enhancement module can adaptively recalibrate channel-wise feature responses by explicitly modeling interdependencies between channels, and reweighting the input feature map to emphasize informative information. Consequently, our network efficiently highlights significant gait-related features while reducing the impact of redundant features, leading to more accurate gait recognition. Lastly, we combine gait-saliency enhancement features of various strips and obtain the final gait feature $f_{\beta}\in\mathbb{R}^{c}$ ($c=256$) for further loss computation through a feature head with multi-FC layers.
5. Experiments
Table 3. Evaluation with different attributes on SUSTech1K valid + test set. We compare our method with silhouette-based SOTA method GaitBase, LiDAR-based SOTA method PointMLP and LiDARGait.
@@ 5. Experiments
Silhouette GaitBase (Fan and etc, 2023) 83.09 79.34 50.95 76.98 77.34 77.31 83.46 26.65 77.50 90.22
Point Cloud PointMLP (Ma et al., 2022) 76.03 71.91 57.09 68.08 58.29 63.28 79.25 70.75 68.86 90.32
LiDARGait (Shen et al., 2023) 91.91 88.61 75.27 88.99 67.55 81.19 94.73 90.04 86.81 95.98
@@
5.1. Evaluation Protocol
@@ 5.2. Implementation Details
For FreeGait, we resize silhouettes to $128\times 88$, similar to Gait3D (Zheng and etc., 2022). Each LiDAR point cloud frame is resampled to $N=256$ points via farthest point sampling(FPS), normalized by setting its center as the origin while preserving original orientations. Range views are cropped and resized to $64\times 64$, following LiDARGait (Shen et al., 2023). For SUSTech1K (Shen et al., 2023), we resample the point clouds to 512 points using FPS, as they have an average of 800 points compared to FreeGait’s 300 points. Other input settings remain consistent with the paper. All methods, except LiDARGait and our HMRNet, are trained with Adam optimizer with a weight decay of 0.0005 and an initial learning rate of 0.001 on the two benchmarks. We employ multi-step learning rate reduction by a factor of 0.1 at the 10,000th and 30,000th iterations, with a total of 50,000 iterations. For LiDARGait and our HMRNet, we maintain the same training procedure on both benchmarks. During the training stage, each input sequence consists of 10 frames, approximating the average length of a human gait cycle. In the training stage, we leverage a combined loss consisting of $BA^{+}$ triplet loss $L_{tri}$ (Chao and etc, 2019) and cross-entropy loss $L_{cls}$ (Lin et al., 2021) with weight parameters $\alpha=1$ and $\beta=0.1$ to train our network. The overall loss used to train the proposed network can be formulated as
(5) $L=\alpha L_{tri}+\beta L_{cls},$
5.3. Comparison with SOTA Methods
@@
Table 2 compares our method with state-of-the-art (SOTA) gait recognition techniques on the FreeGait and SUSTech1K datasets. Our approach outperforms exisiting sota methods LiDARGait (Shen et al., 2023), showing improvements of $\mathbf{6.61\%}$ and $\mathbf{3.42\%}$ in Rank-1 accuracy for FreeGait and SUSTech1K, respectively. These achievements stem from our effective combination of 3D geometry and dynamic motion gait features. Silhouette-based methods suffer from distorted body shapes in view-dependent images, impacting performance. Though 3D SMPL and KP of FreeGait and SUSTech1K are based on the SOTA LiDAR-based mocap method (Zhao et al., 2022), their accuracy is yet to be validated. We evaluated point-based gait recognition using PointNet(Qi and etc, 2017) and PointMLP(Ma et al., 2022) extractors, followed by a recognition head. These point-based methods grapple with sparse representation especially in distant scenes, but PointMLP(Ma et al., 2022), leveraging local feature grouping, showcases the promise in point-based gait recognition. LiDARGait (Shen et al., 2023) mitigates point sparsity and disorder by converting them into a dense range-view representation, achieving notable results. However, this conversion can compromise the original rich 3D geometric and dynamic motion data from raw point clouds. Conversely, our HMRNet extracts dense body structures from range views and accurately captures the raw point clouds’ geometry and motion. It hierarchically processes both local fine-grained and global semantic features, benefiting gait feature extraction.
@@
Figure 6. The exemplar range views on SUSTech1K (Shen et al., 2023) in four different attributes with serious appearance variance between gallery set and probe set.
5.4. Ablation for Network Design.
@@
Table 4. Ablation studies for network modules on FreeGait.
@@
Table 5. Rank-1 accuracy under cross views and low-light conditions on FreeGait dataset.
@@
Silhouette GaitBase (Fan and etc, 2023) 30.31 30.83
Point Cloud LiDARGait (Shen et al., 2023) 59.62 73.33
@@
5.5. More Analysis of Robustness.
@@
Figure 7. Average rank-1 accuracy with increasing testing frames on FreeGait dataset.